OpenStudy (anonymous):

An equilateral triangle has a perimeter of 66 m and a height of 19.1 m. What is its area? I'm so confused.

OpenStudy (daniel14):

hmmm

OpenStudy (anonymous):

Yeah I'm so confused and there aren't any directions

OpenStudy (daniel14):

is there any options?

OpenStudy (anonymous):

Nope :/

OpenStudy (daniel14):

@EclipsedStar

OpenStudy (anonymous):

I'm sooo lost

OpenStudy (daniel14):

dude i feel u ill try to get someone to help u

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

@geerky42 can u help?

geerky42 (geerky42):

"An equilateral triangle has a perimeter of 66 m" So sum of all three sides is 66m. So what is length of one side? WE can treat one side as base.

geerky42 (geerky42):

All sides of equilateral triangle are same, just so you know.

OpenStudy (anonymous):

So it should be 66/3 and that will be one of the sides?

geerky42 (geerky42):

Right?

OpenStudy (anonymous):

Okay so it should be 22, and then multiply that by 19.1 for the area?

OpenStudy (anonymous):

And then the area= 420.2 right?

geerky42 (geerky42):

Actually \(A = \dfrac{1}{2}bh\) So you should divide by 2.

geerky42 (geerky42):

22 multiply by 19.1 then divide by 2.

OpenStudy (anonymous):

210.1?

geerky42 (geerky42):

Then that would be your final answer.

geerky42 (geerky42):

Yep

geerky42 (geerky42):

Does that make sense?

OpenStudy (anonymous):

Yay! Thank you!

OpenStudy (anonymous):

Do you think you could help with another if it isnt a bother?

geerky42 (geerky42):

Sure

OpenStudy (anonymous):

Okay I need to find the area of this polygon:

OpenStudy (anonymous):

I think I know how to do it but I'm not sure

geerky42 (geerky42):

I would ,like you to try it on your own.

geerky42 (geerky42):

Tell me answer and I will check

OpenStudy (anonymous):

The triangle is half of the square, I believe, and the measurements of the triangle are a hypotenuse of 5 and a base of 3, multiply those and then multiply by two to get the whole square?

geerky42 (geerky42):

That "square" isn't really square. It's rectangle. And You would need to find height. Focus on that right triangle part. Use pythagorean theorem to find height.

OpenStudy (anonymous):

Kk. 3^2+b^2=5^2 9+b^2=25 25-9= 16 b=4

geerky42 (geerky42):

Correct, so height is 4 ft. Now can you find area of triangle and rectangle?

OpenStudy (anonymous):

24 for the rectangle 12 for the triangle

geerky42 (geerky42):

Don't forgot that in triangle, you need to divide base times height by 2.

OpenStudy (anonymous):

so it'd be 6?

geerky42 (geerky42):

Yep

geerky42 (geerky42):

So what is total area?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

or is it still 24?

geerky42 (geerky42):

For area of rectangle, you just multiply base by height and that it. For triangle, you also need to divide by 2. Because basically triangle is half the rectangle.